# Orientation tracking method for anisotropic particles

### v1

@@ I Introduction
Tracking objects over time is vital for understanding dynamical systems and behaviours, ranging from pedestrian behaviour in crowds Corbetta et al. (2018), sheep herd behaviour in a bottleneck Garcimartín et al. (2015), to the transportation of particles on the water surface Salmon et al. (2023). In fluid mechanics, particle tracking is a valuable instrument for investigating flows, particle-fluid interactions, and particle dynamics. To most effectively use particle tracking as a measurement tool, the particle tracking method should ideally be able to track multiple (types of) particles simultaneously, should be flexible in camera arrangement (the relative location and orientation of the cameras), and track both the location and orientation of the particles. Preferably the tracking method can track anisotropic particles, since particles in general are anisoptropic, and isotropic particles are the exception. Hitherto scientific research has mainly focused on isotropic (spherical) particles Brown, Warhaft, and Voth (2009); Chiarini and Rosti (2024), though recently some have studied anisotropic particles Voth and Soldati (2017). These particles range from rigid fibres Bakhuis et al. (2017) and spheroids Will et al. (2021), to flexible fibres and discs Verhille and Bartoli (2016); Verhille (2022), hexagonal platelets Tinklenberg, Guala, and Coletti (2023), and particles which break mirror symmetry Piumini et al. (2024). Lagrangian tracking of these particles provides insights into particle dynamics, where both the particle’s location and orientation are relevant for studying the flow-particle and particle-particle interactionsToschi and Bodenschatz (2009); Brandt and Coletti (2022); Harms, Brunton, and McKeon (2024). Many different tracking and measuring methods have been developed for obtaining the location and orientation data of spheres, fibres, and various anisotropic particles Cole et al. (2016); Mathai et al. (2016); Zimmermann et al. (2012); Ouellette, Xu, and Bodenschatz (2006); Ibarra, Bartoli, and Verhille (2023). In this work, we describe a tracking method which finds the location and orientation of anisotropic particles using multiple cameras to image the particles. This method is flexible in terms of the camera setup, and works for particles with concave shapes, even those which have a centre of mass outside of the particle. Previous works have demonstrated 3D orientation tracking methods based on matching multiple 2D orientations Parsa et al. (2012), or finding a particle’s edges Brizzolara et al. (2021). More complex algorithms include methods using a particle marked with a black-and-white pattern Zimmermann et al. (2011); Mathai et al. (2016); Niggel et al. (2023) or projecting a particle model onto the experimental images Marcus et al. (2014); Cole et al. (2016), which are often limited to tracking a single particle. This tracking method distinguishes itself from others by tracking multiple particle (types) simultaneously, being applicable to a wide range of particle shapes and sizes, and works for single or multiple cameras. See Figure 1, the shown particles can be tracked using this method (including particles of convex shapes), and the robustness of the method is carefully examined, such that it can be used as a tool for accurate location and orientation tracking.
@@ II Tracking principles
A wide range of particle tracking algorithms have been developed, with different strong and weak points. Methods based on finding the 2D orientation or detecting the particle endpoints in 2D are very effective for rod-like particles Parsa et al. (2012); Bakhuis et al. (2017); Brizzolara et al. (2021), but are not suited for more complex 3D shapes. A more general approach for 3D particle tracking is carving voxels based on multiple camera images, then finding the particle orientation by fitting the particle to the voxels Alipour, Paoli, and S. Ghaemi (2021); Giurgiu et al. (2024). While this works well for many particles when multiple cameras are set up perpendicular to each other, for our concave particles this method is not sufficient when there is limited optical access (like in a Taylor–Couette setup). To have a flexible camera arrangement to track the anisotropic particles shown in Figure 1, we therefore defer to another method.
@@
We build on these techniques, but now utilise the silhouette of the particle to determine its orientation. We synthetically generate silhouettes of a particle with known orientation and compare to the experimentally recorded silhouette to determine the particle orientation. The method in this paper works in principle for any type of particle which is anisotropic to such a degree that the shape outline can be used to determine its orientation. The working principle can be used for any number of cameras in freely chosen camera arrangements. We expand on previous literature by tracking multiple types of complex-shaped particles simultaneously, and by extensively quantifying the accuracy and robustness for this method, including effects of noise, image resolution and camera arrangements.
@@ II.1 Initialisation
The next step is generating a library of standard orientations, similar to Zimmermann et al. and Mathai et al Zimmermann et al. (2011); Mathai et al. (2016). This library is specific to the particle and the camera parameters, and will later be used to find a first guess for the particle orientation. To generate the library, a standard orientation is defined first, where the vertices of the particle are defined as position vectors . Rotating the particle is done by rotating the vertices using quaternions. To write the particle as a quaternion, the particle vertex vectors are added with a zero:
@@
Following the flowchart of Figure 2, the experimental images are recorded, after which these images are processed, as shown in Figure 5. The left panel shows the raw image from one of the cameras. The particles are extracted from the image as seen in the top-right image. The image is binarised as shown in the middle-right panel of Figure 5.
@@ III.1 Noise
The PDFs of the orientation errors demonstrate that the orientation detection is very accurate overall, with small errors in the detected orientation. The edge noise has a clear effect on the orientation errors, but the method is robust enough to still be very accurate in finding the orientation. The peaks in the PDFs are not located at , which can be explained by the limited resolution of the images. As a result, two very close (but distinct) orientations give the same projections at low resolutions. Therefore, there exists an upper limit for the accuracy of this method, beyond which the reference image and the optimised image cannot be distinguished. The value of this upper limit is dependent on the image resolution, where a higher resolution image leads to lower errors at the accuracy limit, as illustrated in the section below.
@@ III.4 Camera arrangements
We find that using more cameras is in general more accurate, noting that the way the cameras are arranged is also important. The PDFs show that the tetrahedral setup of cameras is more accurate than the near-planar camera setup.
The near-planar camera arrangement still performs slightly better than the three orthogonal cameras, indicating that using more cameras is in general beneficial, even when these cameras are arranged in a near-planar manner. Using a single camera, we notice that the value for is larger than for other arrangements. This is caused by the algorithm finding a number of incorrect orientations, since the particle orientation cannot be determined for specific orientations, resulting in large values of ). Still, more than of the orientations are found with a small error (). This shows that the orientation finding method still works in principle, but cannot be reliably applied to this particle for a single camera.
@@ Acknowledgements.
The authors thank Detlef Lohse and Federico Toschi for their insights and discussions. We would like to thank Gert-Wim Bruggert, Thomas Zijlstra, and Martin Bos for their technical support. This research has been funded by the Dutch Research Council (NWO) under grant OCENW.GROOT.2019.031. Elian Bernard visited the Physics of Fluids department under an Erasmus grant. The authors report no conflict of interest.References
- Corbetta et al. (2018) A. Corbetta, J. A. Meeusen, C. Lee, R. Benzi, and F. Toschi, Phys. Rev. E 98 (2018).
- Garcimartín et al. (2015) A. Garcimartín, J. M. Pastor, L. M. Ferrer, J. J. Ramos, C. Martín-Gómez, and I. Zuriguel, Phys. Rev. E 91 (2015).
- Salmon et al. (2023) H. R. S. Salmon, L. J. Baker, J. L. Kozarek, and F. Coletti, Water Resour. Res. 59 (2023).
- Brown, Warhaft, and Voth (2009) R. D. Brown, Z. Warhaft, and G. A. Voth, Phys. Rev. Lett. 103 (2009).
- Chiarini and Rosti (2024) A. Chiarini and M. Rosti, J. Fluid Mech. 988 (2024).
- Voth and Soldati (2017) G. A. Voth and A. Soldati, Annu. Rev. Fluid Mech. 49 (2017).
- Bakhuis et al. (2017) D. Bakhuis, V. Mathai, R. A. Verschoof, R. Ezeta, D. Lohse, S. G. Huisman, and C. Sun, Annu. Rev. Fluid Mech. 49 (2017).
- Will et al. (2021) J. B. Will, V. Mathai, S. G. Huisman, D. Lohse, C. Sun, and D. Krug, J. Fluid Mech 912 (2021).
- Verhille and Bartoli (2016) G. Verhille and A. Bartoli, Exp. Fluids 57 (2016).
- Verhille (2022) G. Verhille, J. Fluid Mech. 933 (2022).
- Tinklenberg, Guala, and Coletti (2023) A. Tinklenberg, M. Guala, and F. Coletti, J. Fluid Mech. 962 (2023).
- Piumini et al. (2024) G. Piumini, M. P. A. Assen, D. Lohse, and R. Verzicco, J. Fluid Mech. 995 (2024).
- Toschi and Bodenschatz (2009) F. Toschi and E. Bodenschatz, Annu. Rev. Fluid Mech. 41 (2009).
- Brandt and Coletti (2022) L. Brandt and F. Coletti, Annu. Rev. Fluid Mech. 54 (2022).
- Harms, Brunton, and McKeon (2024) T. Harms, S. L. Brunton, and B. J. McKeon, ArXiv (2024).
- Cole et al. (2016) B. C. Cole, G. G. Marcus, S. Parsa, S. Kramel, R. Ni, and G. A. Voth, J. Vis. Exp. 112 (2016).
- Mathai et al. (2016) V. Mathai, M. W. M. Neut, E. P. van der Poel, and C. Sun, Exp Fluids 57 (2016).
- Zimmermann et al. (2012) R. Zimmermann, L. Fiabane, Y. Gasteuil, R. Volk, and J.-F. Pinton, Phys. Scr. (2012).
- Ouellette, Xu, and Bodenschatz (2006) N. T. Ouellette, H. Xu, and E. Bodenschatz, Exp. Fluids 40 (2006).
- Ibarra, Bartoli, and Verhille (2023) E. Ibarra, A. Bartoli, and G. Verhille, Exp. Fluids 64 (2023).
- Parsa et al. (2012) S. Parsa, E. Calzavarini, F. Toschi, and G. A. Voth, Phys. Rev. Lett. 109 (2012).
- Brizzolara et al. (2021) S. Brizzolara, M. E. Rosti, S. Olivieri, L. Brandt, M. Holzner, and A. Mazzino, Phys. Rev. X 11 (2021).
- Zimmermann et al. (2011) R. Zimmermann, Y. Gasteuil, M. Bourgoin, R. Volk, A. Pumir, and J. Pinton, Rev. Sci. Instrum. 82 (2011).
- Niggel et al. (2023) V. Niggel, M. R. Bailey, C. van Baalen, N. Zosso, and L. Isa, Soft Matter 19 (2023).
- Marcus et al. (2014) G. G. Marcus, S. Parsa, S. Kramel, R. Ni, and G. A. Voth, New J. Phys. 16 (2014).
- Alipour, Paoli, and S. Ghaemi (2021) M. Alipour, M. D. Paoli, and A. S. S. Ghaemi, J. Fluid Mech. 916 (2021).
- Giurgiu et al. (2024) V. Giurgiu, G. C. A. Caridi, M. D. Paoli, and A. Soldati, Phys. Rev. Lett. 133 (2024).
- Tsai (1987) R. Tsai, IEEE J. Robot. Autom. 3 (1987).
- Bourgoin and Huisman (2020) M. Bourgoin and S. G. Huisman, Rev. Sci. Instrum. 91 (2020).
- Nelder and Mead (1965) J. A. Nelder and R. Mead, Comput. J (1965).
- van Gils et al. (2011) D. P. M. van Gils, G. W. H. Bruggert, D. P. Lathrop, C. Sun, and D. Lohse, Rev. Sci. Instrum. 82 (2011).
- Huisman et al. (2015) S. G. Huisman, R. C. A. van der Veen, G. W. H. Bruggert, D. Lohse, and C. Sun., Rev. Sci. Instrum. 86 (2015).
- van der Veen et al. (2016) R. C. A. van der Veen, S. G. Huisman, O. Y. Dung, H. L. Tang, C. Sun, and D. Lohse, Phys. Rev. Fluids 1 (2016).
- Porta, Voth, and Crawford (2001) A. L. Porta, G. A. Voth, and A. Crawford, Nature 40 (2001).
- Calzavarini et al. (2009) E. Calzavarini, R. Volk, M. Bourgoin, E. Lévêque, J.-F. Pinton, and F. Toschi, J. Fluid Mech. 630 (2009).
- Klein et al. (2012) S. Klein, M. Gibert, A. Bérut, and E. Bodenschatz, Meas. Sci. Technol. 24 (2012).
- Zimmermann et al. (2010) R. Zimmermann, H. Xu, Y. Gasteuil, M. Bourgoin, R. Volk, J.-F. Pinton, and E. Bodenschatz, Rev. Sci. Instrum. 81 (2010).

### v2

# Multi-camera orientation tracking method for anisotropic particles in particle-laden flows

@@ I Introduction
Tracking objects over time is vital for understanding dynamical systems and behaviours, ranging from pedestrian behaviour in crowds Corbetta et al. (2018), sheep herd behaviour in a bottleneck Garcimartín et al. (2015), to the transportation of particles on the water surface Salmon et al. (2023). In fluid mechanics, particle tracking is a valuable instrument for investigating flows, particle-fluid interactions, and particle dynamicsSchröder and Schanz (2022). Tracking small tracer particles can be used to visualise complex dynamics in flows Godbersen et al. (2021). Other fluid-dynamics applications of particle tracking are studying particle clustering dynamics Petersen, Baker, and Coletti (2019), investigating the settling of snowflakes Nemes et al. (2017), or measuring the dynamics of complex particle geometries Collins et al. (2021). To most effectively use particle tracking as a measurement tool, the particle tracking method should ideally be able to track multiple (types of) particles simultaneously, should be flexible in camera arrangement (the relative location and orientation of the cameras), and track both the location and orientation of the particles. Preferably the tracking method can track anisotropic particles, since particles in general are anisoptropic, and isotropic particles are the exception. Hitherto scientific research has mainly focused on isotropic (spherical) particles Brown, Warhaft, and Voth (2009); Chiarini and Rosti (2024), though recently some have studied anisotropic particles Voth and Soldati (2017). These particles range from rigid fibres Bakhuis et al. (2019) and spheroids Will et al. (2021), to flexible fibres and discs Verhille and Bartoli (2016); Verhille (2022), hexagonal platelets Tinklenberg, Guala, and Coletti (2023), and particles which break mirror symmetry Piumini et al. (2024). Lagrangian tracking of these particles provides insights into particle dynamics, where both the particle’s location and orientation are relevant for studying the flow-particle and particle-particle interactionsToschi and Bodenschatz (2009); Brandt and Coletti (2022); Harms, Brunton, and McKeon (2024). Many different tracking and measuring methods have been developed for obtaining the location and orientation data of spheres, fibres, and various anisotropic particles Cole et al. (2016); Mathai et al. (2016); Zimmermann et al. (2012); Ouellette, Xu, and Bodenschatz (2006); Ibarra, Bartoli, and Verhille (2023). In this work, we describe a tracking method which finds the location and orientation of anisotropic particles using multiple cameras to image the particles. This method is flexible in terms of the camera setup, and works for particles with concave shapes, even those which have a centre of mass outside of the particle. Previous works have demonstrated 3D orientation tracking methods based on matching multiple 2D orientations Parsa et al. (2012), finding a particle’s edges Brizzolara et al. (2021), or a combination of the two Verhille and Bartoli (2016), where multiple particles can be tracked simultaneously Bounoua, Bouchet, and Verhille (2018); Shaik et al. (2020); Baker and Coletti (2022). More complex algorithms include methods using a particle marked with a black-and-white pattern Zimmermann et al. (2011); Mathai et al. (2016); Niggel et al. (2023) or projecting a synthetic particle model onto the experimental images Marcus et al. (2014); Cole et al. (2016). This tracking method distinguishes itself from others by tracking multiple particle types simultaneously, being applicable to a wide range of particle shapes and sizes, including those that are concave and have their centre of mass outside the body. In addition, this method works for single or multiple cameras, and the camera arrangement is flexible. See Figure 1, the shown particles can be tracked using this method (including particles of convex shapes), and the robustness of the method is carefully examined, such that it can be used as a tool for accurate location and orientation tracking.
@@ II Tracking principles
A wide range of particle tracking algorithms have been developed, with different strong and weak points. Methods based on finding the 2D orientation or detecting the particle endpoints in 2D are very effective for rod-like particles Parsa et al. (2012); Bakhuis et al. (2019); Brizzolara et al. (2021), but are not suited for more complex 3D shapes. A more general approach for 3D particle tracking is carving voxels based on multiple camera images, then finding the particle orientation by fitting the particle to the voxels Alipour, Paoli, and S. Ghaemi (2021); Giurgiu et al. (2024). This method is shown to work well for tracking, and allows for a wide variety of geometries or configurations. Fu and Liu (2018); Masuk, Salibindla, and Ni (2019). However, for our chiral particles, for specific camera configurations (for example when the cameras are closely arranged in a near-planar arrangement), we cannot distinguish the handedness of chiral the particle using this carving method. Therefore, for the chosen particle geometries, we do not have full camera flexibility when using the carving method. To have a flexible camera arrangement to track the anisotropic particles shown in Figure 1, we therefore defer to another method.
@@
We build on these techniques, but now utilise the silhouette of the particle to determine its orientation. We synthetically generate silhouettes of a particle with known orientation and compare to the experimentally recorded silhouette to determine the particle orientation. The method in this paper works in principle for any type of particle which is anisotropic to such a degree that the shape outline can be used to determine its orientation. The working principle can be used for any number of cameras in freely chosen camera arrangements. The materials needed to implement this method is therefore limited to high-speed cameras (two being the minimum for complex particles, and each extra camera improves the accuracy of the method) and a computer for processing the recorded images. We expand on previous literature by tracking multiple types of complex-shaped particles simultaneously, and by extensively quantifying the accuracy and robustness for this method, including effects of noise, image resolution and camera arrangements.
@@ II.1 Initialisation
The next step is generating a library of standard orientations, similar to Zimmermann et al. and Mathai et al Zimmermann et al. (2011); Mathai et al. (2016). This library is specific to the particle and the camera parameters, and will later be used to find a first guess for the particle orientation. Although this first guess could also be found by using a deep learning model to give an orientation as recently shown by di Giusto et al.Giusto et al. (2024), who used a neural network to determine the particle orientation based on images of particle projections. To generate the library, a standard orientation is defined first, where the vertices of the particle are defined as position vectors . Rotating the particle is done by rotating the vertices using quaternions. To write the particle as a quaternion, the particle vertex vectors are added with a zero:
@@
Following the flowchart of Figure 2, the experimental images are recorded, after which the particles are isolated and extracted. We remove any small bubbles, as only particles should be extracted from the images. This can be achieved using a segmentation method of choice Gonzalez and Woods (2002); Rother, Kolmogorov, and Blake (2004); Bruijne and Nielsen (2004) where necessary. Overlapping particles are cut out as a single image, but cannot be distinguished, and the overlapping particles’ orientations cannot be easily found in the current state of the described algorithm. In theory, one could optimise the orientation of two particles simultaneously, but the large amount of local minima makes this a challenging endeavour. Once the particle images are extracted from the raw image, these images are processed, as shown in Figure 5. The left panel shows the raw image from one of the cameras. The particles are extracted from the image as seen in the top-right image. The image is binarised as shown in the middle-right panel of Figure 5.
@@ II.2 Determining the orientation
As mentioned before, particles which overlap on one of the cameras could not be reliably distinguished and tracked while overlapping, which is a limitation in the current algorithm, and limits its use to cases of dilute particle concentrations. This limitation may be overcome by using the known particle shape for specific geometries Shen et al. (2000); Neoptolemou et al. (2022), or a machine-learning aided approach may be used to better distinguish the overlapping particles Puzyrev et al. (2020).
@@ II.3 Correcting the centre of mass
If a high degree of accuracy is required, the corrected 3D position can be used to generate the camera vectors for the orientation optimisation, thereby iteratively optimising the position and orientation of the tracked particle which quickly converges. Alternatively, a simultaneous optimisation of the position and orientation could be implemented, as shown in other work Oehmke et al. (2021).
@@ III Robustness & errors
This section only considers the error in the orientation, since this method focuses on finding the orientation. How the error in position affects the orientation and how the orientation error affects the corrected is shown in Appendix A. Furthermore, only the errors of the orientation of the chiral particles are investigated here. The results for all investigated particle geometries are comparable, as shown in Appendix B.
@@ III.1 Noise
The PDFs of the orientation errors demonstrate that the orientation detection is very accurate overall, with small errors in the detected orientation. Here we note that the orientation error depends on how ’jagged’ the particle silhouette is; more edges and corners lead to a smaller orientation error, whereas a more ’smooth’ particle silhouette gives a slightly larger error. The edge noise has a clear effect on the orientation errors, but the method is robust enough to still be very accurate in finding the orientation. The mean orientation errors for this synthetic dataset using our tracking method are of order or less, which is comparable to or smaller than the errors in similar orientation tracking methods Zimmermann et al. (2011); Mathai et al. (2016). The peaks in the orientation error PDFs are not located at , which can be explained by the limited resolution of the images. As a result, two very close (but distinct) orientations give the same projections at low resolutions. Therefore, there exists an upper limit for the accuracy of this method, beyond which the reference image and the optimised image cannot be distinguished. The value of this upper limit is dependent on the image resolution, where a higher resolution image leads to lower errors at the accuracy limit, as illustrated in the section below.
@@ III.4 Camera arrangements
We find that using more cameras is in general more accurate, noting that the way the cameras are arranged is also important. The PDFs show that the tetrahedral setup of cameras is more accurate than the near-planar camera setup. Therefore, we find that the way the cameras are arranged also affects the orientation error, but changing the number of used cameras has a greater effect. This is evident when comparing the near-planar arrangement (which uses 4 cameras) to the three orthogonal cameras, where the near-planar arrangement has a smaller average orientation error compared to the three orthogonal cameras.
Using a single camera, we notice that the value for is larger than for other arrangements. This is caused by the algorithm finding a number of incorrect orientations, since the particle orientation cannot be determined for specific orientations, resulting in large values of ). Still, more than of the orientations are found with a small error (). This shows that the orientation finding method still works in principle, but cannot be reliably applied to this particle for a single camera.
@@ Acknowledgements.
The authors thank Detlef Lohse and Federico Toschi for their insights and discussions. We would like to thank Gert-Wim Bruggert, Thomas Zijlstra, and Martin Bos for their technical support. This research has been funded by the Dutch Research Council (NWO) under grant OCENW.GROOT.2019.031. Elian Bernard visited the Physics of Fluids department under an Erasmus grant. The authors report no conflict of interest.Appendix A Error propagation between position and orientation
The described algorithm treats the finding of the position and the finding of the orientation separately. However, the error in the position slightly affects the orientation, in the sense that the position determines the camera vector along which the synthetic particle model is projected.
For the experiment tracking many chiral particles, as shown in section IV, the average image size is approximately . An error in the position of of the image size (which is approximately the biggest distance between the projected centre of mass, and the centroid of the projection) leads to a difference in camera vector of approximately . This projection error leads to an orientation error of similar magnitude, which is comparable to the mean orientation error of the optimisation process.
This error is sufficiently small for the fluid-dynamics applications this method was developed for. If a higher accuracy is required, more iterations of the algorithm can be used, or the location and orientation can be optimised simultaneously.
Since the orientation is used to correct the centre of mass position, the error in orientation affects the error in position. We test the effect of the orientation error by calculating the displacement of the projected centre of mass (in pixels), as a result of an erroneous orientation.
A set of 1000 random orientations is generated, where each orientation is randomly rotated 100 times by the orientation error. This calculation is done for three different square image sizes: , , and , where we use the corresponding mean orientation error as found in Figure 9 to rotate the reference orientations. The displacement of the projected centre of mass (in pixels) is then calculated for each camera. The displacements are averaged per image size, and calculated as a percentage of the image size: these results are shown in Figure 21.
This figure shows that the projected centre of mass is displaced by less than of the image size.
For the shown chiral particle setup, this gives a position error of () resulting from the orientation error. The magnitude of this error is sufficiently small for the fluid-dynamics applications for which this algorithm was developed.
Appendix B Orientation errors for different particles
So far, the robustness and the orientation errors have only been evaluated for the chiral particles. Here we evaluate the orientation errors for the other shapes in this manuscript (the tetrad and oloid), for different image sizes. The synthetic data is generated as before, now using 1000 orientations to determine the mean orientation error. The mean orientation errors for all used shapes are shown in Figure 22, for multiple image sizes.
This plot shows that the magnitude of the mean orientation error differs slightly per particle shape, but the overall scaling with image size is similar for all shapes. The mean orientation errors are sufficiently small for our applications for all particle shapes.